OpenStudy (anonymous):

Which Equation Correctly Demonstrates The Distributive Property A. -3(4-2x)=-12+6x B. -3(4-2x)=-12-2x C. -3(4-2x)=12-6x D. -3(4-2x)=-12+2x

jagr2713 (jagr2713):

Ok do you know what the distributive property is?

OpenStudy (anonymous):

Not At All.

jagr2713 (jagr2713):

Ok its like this Imagine we have 3(2x+6) distributive property is simply multiplying 3 to all the numbers inside the parentheses so we get 

jagr2713 (jagr2713):

This explains more @Imperfect_Queen http://www.coolmath.com/prealgebra/06-properties/05-properties-distributive-01 So from all that what do you think the answer is?

OpenStudy (anonymous):

C. -3(4-2x)=12-6x

jagr2713 (jagr2713):

You forgot something. This is a costly mistake whats is -3*4

OpenStudy (anonymous):

-12

jagr2713 (jagr2713):

Good so now try again :D

OpenStudy (anonymous):

A. -3(4-2x)=-12+6x

OpenStudy (anonymous):

@jagr2713

jagr2713 (jagr2713):

Correct great job :D Do you fully understand distributive property?